one of the factors is chosen at random. what is the probability that the chosen factor is not a multiple of 4?
Answer by ewatrrr(24785)

Hi,
An ex where 24 was the number given and factors listed:
the 8 factors of 24 are: 1, 2, 3, 4, 6, 8, 12, 24
P(factor chosen is not a multiple of 4) = 4/8 = 1/2